OpenStudy (anonymous):

Suppose the temperature increases 8° to -7°F what was the starting temperature A.15° B -20° C.-13° D.-15°

OpenStudy (anonymous):

8+7=15 but because you are adding to get -7, you will have -15 (D) for your answer.

OpenStudy (phi):

If you think like they teach in Algebra x = unknown start temperature (unknown) temperature increases 8° write that as x + 8 (8 more than the unknown temperature x) to -7°F after adding 8 (increasing by 8) we get to -7 x+8 = -7 once we have an equation, we can solve it. add -8 to both sides x+8 -8 = -7 -8 on the left side 8-8 is 0 . on the right side -7-8= -15 x+0 = -15 or simply x = -15 the unknown temperature is -15

OpenStudy (skullpatrol):